OpenStudy (anonymous):

Find the surface area of a prism with bases which are equilateral triangles of side length 4 cm and a height of 5 cm.

OpenStudy (raden):

surface area = base area + (perimeter of base * the height of that prism)

jhonyy9 (jhonyy9):

try drawing it

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

area of base can you calculing ?

jhonyy9 (jhonyy9):

what is formula for area of a triangle ?

jhonyy9 (jhonyy9):

are you here ?

OpenStudy (anonymous):

A=1/2bh the area of the base of the prism is 6.93

jhonyy9 (jhonyy9):

ok and area of a lateral side =?

jhonyy9 (jhonyy9):

but can rewriting here how you have got 6,93 for area of base ?

OpenStudy (anonymous):

the base of the prism is a triangle so I used the formula h=\[(\sqrt{3}*s)/2\] s being 4 and the answer I came up with for the height was 3.46 and then i plugged that in to the equation i used to find the area of the base

jhonyy9 (jhonyy9):

ok than now you ned to calculi arae of a lateral side = ?

jhonyy9 (jhonyy9):

or us formula what have wrote firstly @RadEn

OpenStudy (anonymous):

ok if i use @RadEn's formula my answer is 66.92

jhonyy9 (jhonyy9):

write please here your calculi

OpenStudy (anonymous):

6.92+(12*5) =6.92+60 =66.92

jhonyy9 (jhonyy9):

ok do you understand it now sure right ?

OpenStudy (anonymous):

yes, thank you!

jhonyy9 (jhonyy9):

yw good luck bye